OpenStudy (anonymous):

Simplify (x^-2/5y^1/3)^15 x^6y^5 x^6/y^5 y^5/x^6 1/x^6y^5

OpenStudy (jhannybean):

\[\huge \left( \frac{\frac{x^{-2}}{5y^1}}{3}\right)^{15}\] Is this it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(x^2y)^-15

OpenStudy (anonymous):

which means (1) / [(x^30)(y^15)]

OpenStudy (anonymous):

u got it?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

u need th others simplified too?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh my im so sorry about this.. I had written the question incorrectly the actual answer is [1/15x^2y]^15

OpenStudy (anonymous):

\[[\frac{ 1 }{ 15x^2y }]^(15)\]

OpenStudy (jhannybean):

\[\huge \left( \frac{\frac{x^{-2}}{5y^1}}{3}\right)^{15}\] First simplify the stuff inside the parenthesis \[\large \left(\frac{x^{-2}}{15y}\right)^{15}\]distribute the 15 to everything in the numerator and denominator \[\large \left( \frac{x^{(-2\cdot 15)}}{(15 \cdot 15)y^{15}}\right)\] simplify \[\large \left( \frac{x^{-30}}{225\cdot y^{15}}\right)\] \[\large\left( \frac{1}{225 \cdot x^{30} \cdot y^{15}}\right)\]

OpenStudy (jhannybean):

@chrissy401 is what you've written to the 15th power?

OpenStudy (anonymous):

i wrote the entire thing to the 15th power so as not to confuse the equation

OpenStudy (anonymous):

If i brought in the power.. We would get the same answer

OpenStudy (jhannybean):

\[\large \left(\frac{1}{15x^2y}\right)^{15} = \left( \frac{1}{(15^{15})(x^{30})(y^{15})}\right)\]\[\large 15^{15} \neq (15\cdot 15)\] I solved it the wrong way as well.

OpenStudy (anonymous):

yeah.. this is why i never bring in powers.. i like leaving them in the raw form.. I would have preferred to leave it like this \[\left[ 15x^2y \right]^(-15)\]